# Effort: Efficient Orthogonal Modeling for Generalizable AI-Generated Image Detection

### v1

@@ Abstract
Existing AI-generated image (AIGI) detection methods often suffer from limited generalization performance. In this paper, we identify a crucial yet previously overlooked asymmetry phenomenon in AIGI detection: during training, models tend to quickly overfit to specific fake patterns in the training set, while other information is not adequately captured, leading to poor generalization when faced with new fake methods. A key insight is to incorporate the rich semantic knowledge embedded within large-scale vision foundation models (VFMs) to expand the previous discriminative space (based on forgery patterns only), such that the discrimination is decided by both forgery and semantic cues, thereby reducing the overfitting to specific forgery patterns. A straightforward solution is to fully fine-tune VFMs, but it risks distorting the well-learned semantic knowledge, pushing the model back toward overfitting. To this end, we design a novel approach called Effort: Efficient orthogonal modeling for generalizable AIGI detection. Specifically, we employ Singular Value Decomposition (SVD) to construct the orthogonal semantic and forgery subspaces. By freezing the principal components and adapting the residual components ($\sim$0.19M parameters), we preserve the original semantic subspace and use its orthogonal subspace for learning forgeries. Extensive experiments on AIGI detection benchmarks demonstrate the superior effectiveness of our approach.
Equal Contribution. Corresponding Author.
Figure 1: Evidence for illustrating the asymmetry phenomenon in AI-generated image detection. We show that the baseline detector (i.e., Xception [70]) tends to quickly overfit to the fake patterns in the training set, causing the generalization issue when facing previously unseen fake methods.
@@ 1 Introduction
The rapid development of AI generative technologies has significantly lowered the barrier for creating highly realistic fake images. As deep generative models advance and mature [21, 28, 68], the proliferation of AI-generated images (AIGIs) has drawn considerable attention, driven by their ability to produce high-quality content with relative ease. However, these advancements also introduce significant risks, if misused for malicious purposes such as deepfakes (mainly including face-swapping [39] and face-reenactment [81]), which may violate personal privacy, spread misinformation, and erode trust in digital media. Consequently, there is an urgent need to develop a reliable and robust framework for detecting AIGIs. [^1] In the context of this research, AIGI primarily refers to deepfakes (face-swapping) and synthetic images (e.g., nature or arts).
Most existing studies in AIGI detection [86, 70] typically approach the real/fake classification problem as a symmetric binary classification task, akin to the “cat versus dog” problem. A standard binary classifier, often based on deep neural networks, is trained to distinguish between real and fake images by predicting the likelihood of a given test image being fake during inference. Although this paradigm yields promising results when the training and testing distributions (in terms of fake generation methods) are similar, its performance tends to degrade significantly when applied to previously unseen fake methods.
Figure 2: Illustration of the comparison without and with incorporating semantic knowledge. Naively trained models rely solely on forgery patterns for discrimination (the left), while incorporating semantic knowledge expands the discrimination space (the right), such that the discrimination is decided by both forgery and semantic cues, thereby alleviating the overfitting to seen fakes.
To understand the underlying reasons for the failure in generalization, we conducted extensive preliminary investigations and identified an asymmetry phenomenon in AIGI detection: naively trained models tend to take the shortcut and quickly overfit the specific fake patterns in the training set, resulting in the undesirable generalization performance when facing novel fake methods. Visualization in Fig. 1 corroborates this observation. Specifically, the vanilla detector (i.e., Xception [70]) quickly fits the fake patterns at the very early training stage (only a few iterations), resulting in a very low loss of fake, while the real loss is significantly higher than the fake loss ($\sim$ 100$\times$ larger). This is likely because existing AIGI detection datasets [70, 86] typically contain limited and homogeneous fake types, while real samples exhibit significantly greater diversity and variance (such as different categories and scenarios). This asymmetry causes the detection model to quickly overfit to the seen forgery patterns during training, limiting its generalization to detect unseen fakes.
To address this, a key insight is to incorporate the rich semantic knowledge embedded within the large-scale vision foundation models (VFMs) to expand the previous discriminative space (relying on forgery pattern only), such that the discrimination is decided by both forgery and semantic cues. We provide an illustration example in Fig. 2 for intuitively understanding the role of semantic knowledge. Specifically, naively trained models rely solely on forgery patterns for discrimination (the left), fitting well to the seen fakes but failing to generalize previously unseen fakes. When semantic knowledge is incorporated, the discrimination space is largely expanded, with different semantic categories such as Bird and Fish separated (the right). In this case, the discrimination is determined by both forgery and semantic cues, thus reducing the overfitting to the seen forgery patterns and enhancing the model’s generalization.
So, is naively fine-tuning a VFM sufficient to obtain a generalizable AIGI detector? Unfortunately, it risks distorting the well-learned semantic knowledge of the original VFM, pushing the VFM model back toward overfitting. The t-SNE results in Fig. 3 show that fully fine-tuning a VFM can still overfit the seen fake patterns, due to the mentioned asymmetry problem. Specifically, both the Xception (Vanilla CNN) and CLIP [65] (VFM) detector group only the specific fake patterns within the training set into a single cluster, while all other data, including real and fake from unseen manipulation methods are classified into the other cluster. This highlights that simply fine-tuning the VFMs (even using CLIP) is still insufficient to overcome the asymmetry issue and achieve robust generalization.
@@
To this end, we design a novel approach called Effort: Efficient orthogonal modeling for generalizable AIGI detection. Specifically, we employ Singular Value Decomposition (SVD) to construct the orthogonal semantic and forgery subspaces. By freezing the principal components and adapting the residual components ($\sim$0.19M parameters), we preserve the original semantic subspace and use its orthogonal subspace for learning forgeries. We conduct experiments on both deepfake detection and synthetic image detection benchmarks and surprisingly find that our approach achieves significant superiority over other SOTAs with very few trainable parameters.
Overall, this work makes the following key contributions:
@@
  • Asymmetry Phenomenon in AIGI Detection: We introduce the concept of asymmetry in AIGI detection, which has often been overlooked in previous works. A naively trained detector tends to quickly fit the seen fake methods well but, in doing so, it often overfits to specific fake patterns in the training set, limiting its generalization ability to detect unseen fake methods.
  • Semantic Knowledge Involved and Orthogonality: We propose a novel approach, Effort, to address the asymmetry in AIGI detection by (1) leveraging the rich semantic knowledge within the Vision Foundation Models (VFMs) to explain the previous discrimination space, alleviating the over-relying on forgery patterns only, and (2) Using SVD to construct the orthogonal semantic and forgery subspaces, avoiding the distortion of original well-learned semantic knowledge during learning fakes.
@@
  • Stricking Result and Efficiency: Effort achieves remarkably high generalization performance on previously unseen forgeries using nearly only 0.19M tunable parameters, making it efficient and potentially scalable. Moreover, our approach can be applied to detect a broad range of AIGIs, including but not limited to face forgeries (e.g., classical deepfakes) and synthetic natural images.
@@ 2 Related Work
Our work focuses on detecting AI-generated images (AIGIs), especially deepfake images (e.g., face-swapping) and synthetic images (e.g., nature or art), following [94]. As the majority of recent works specifically focus on dealing with the generalization issue, where the training and testing distribution differ (in terms of fake methods), we will briefly introduce the classical and recent detection methods toward generalization in deepfake and synthetic images, respectively.
@@
The task of deepfake detection grapples profoundly with the issue of generalization. To tackle the generalization issue, one mainstream approach is fake pattern learning. Most existing works are within this line. These methods generally design a “transformation function”, e.g., frequency transformation [42, 53, 47], blending operations [44, 99, 72, 8], reconstruction [104, 4], content/ID disentanglement [91, 30, 20], to transform the original input $x$ into $x^{\prime}$, where they believe that the more general fake patterns can be captured within the feature space of $x^{\prime}$ compared to $x$. However, given the ever-increasing diversity of forgery methods in the real world, it is unrealistic to elaborate all possible fake patterns and “expect” good generalization on unseen fake methods. Another notable direction is to real distribution learning, with a specific methodology involved: one-class anomaly detection [35, 40]. Specifically, [35] introduced a one-class-based anomaly detection, where “abnormal” data is detected by the proposed reconstruction error as the anomaly score. [40] proposed a similar approach to create pseudo-fake “anomaly” samples by using image-level blending on different facial regions. However, it is challenging to ensure that the detector can learn a robust representation of real images by using the very limited real data in existing deepfake datasets (e.g., the FF++ dataset [70] contains only 1,000 real videos with imbalanced facial attribute distributions [83]).
@@
With the rapid advancement of existing AI generative technologies, the scope of forged content has expanded beyond facial forgeries to encompass a wide range of scenes. In this context, similar to the deepfake detection field, most existing works typically focus on fake pattern learning that mines the low-level forgery clues from different aspects. Specifically, several approaches have been proposed to capture low-level artifacts, including RGB data augmentations [86], frequency-based features [31], gradients [76], reconstruction artifacts [87, 6, 54], and neighboring pixel relationships [79], random-mapping feature [77]. To illustrate, BiHPF [31] amplifies artifact magnitudes through the application of dual high-pass filters, while LGrad [76] uses gradient information from pre-trained models as artifact representations. NPR [79] introduces a straightforward yet effective artifact representation by rethinking up-sampling operations. In addition to learning from scratch, there are also several research works [61, 89, 48] that perform fake pattern learning by leveraging the existing vision foundation models. For instance, UniFD [61] directly freezes the visual encoder of the pre-trained CLIP model and tunes only a linear layer for binary classification, demonstrating effective deepfake detection even with previously unseen sources. LASTED [89] proposes designing textual labels to supervise the CLIP vision model through image-text contrastive learning, advancing the field of synthetic image detection. These arts have shown notable improvement in generalization performance when facing previously unseen fake methods.
Figure 4: The proposed SVD-based adaptation approach for AIGI detection. The left branch is the decomposition matrix of the principle components approximation using SVD, while the right residual branch enables the orthogonal learning of real/fake discriminative features.
@@ 3 Methodology
3.1 Motivation
Detecting AI-generated images (AIGI) has become an increasingly important topic. Existing works typically formulate AIGI detection as a standard binary classification problem (just like “cat versus dog”) and often get unsatisfied generalization performance. Our preliminary investigation surprisingly found that the failure of generalization can be attributed to the previously overlooked asymmetry phenomenon, where a naive trained detector tends to quickly overfit the specific fake pattern present in the training set, thereby limiting its generalization when facing previously unseen fake methods. To better understand the reasons behind the failure generalization, we present three important research questions (RQs) below.
RQ-1: Why do naively trained detectors fail to generalize to new unseen fake methods?
In the previous sections, we have shown that the asymmetry phenomenon, where the naively trained models tend to quickly overfit the specific fake patterns in the training set, resulting in undesirable generalization performance when facing unseen fake methods (see Fig. 1). This is likely because existing AIGI detection datasets [70, 86] typically contain limited and homogeneous fake types, while real samples exhibit significantly greater diversity and variance. This asymmetry causes the detection model to quickly overfit to the specific and limited forgery methods during training, dominating the model’s discrimination space and shaping the whole space to be low-ranked (see Fig. 3(a)), thereby limiting its representational capacity and eventually hurting the generalization. We will provide a deeper analysis and evidence of this claim in Sec. 4.3.
RQ-2: Why incorporating semantic knowledge helps to alleviate the model’s overfitting?
Given RQ-1, to address the asymmetric problem, a key insight is to leverage the rich semantic knowledge of VFMs to expand and enrich the low-ranked discrimination space (relying on forgery patterns only) to higher-ranked, encouraging the final discrimination based on both forgery patterns and semantic knowledge, thereby improving the alleviating the overfitting to the forgery patterns. Recall that there are numerous well-trained and generalizable vision foundation models (VFMs) available such as CLIP [65], SigLIP [96], and BEIT-v2 [63], which have already learned rich and robust semantic representation. Thus, leveraging the extensive semantic knowledge within VFMs presents a reasonable solution to expand the low-ranked discrimination space and reduce overfitting.
RQ-3: Why do naively fine-tuning VFMs fail to generalize to new unseen fake methods?
A straightforward solution that incorporates semantic knowledge into detection is fine-tuning VFMs, but it risks distorting the well-learned semantic knowledge, falling back to the overfitting to the forgery patterns, similar to a vanilla CNN baseline, as shown in Fig. 3. To overcome these issues, we propose to achieve the orthogonality of semantic and forgery subspaces, where the forgery pattern learning is conducted within a new subspace that is orthogonal to the original semantic subspace, avoiding the distortion of the pre-existing semantic knowledge. To achieve this, we introduce a novel approach that uses SVD to explicitly construct the orthogonal semantic and forgery subspaces by freezing the principal components for representing semantic knowledge, while adapting only the residual components for forgery pattern learning.
3.2 Effort Approach
The overall pipeline of the proposed Effort approach is illustrated in Fig. 4, aiming to address the asymmetry phenomena in AIGI detection. In general, our approach involves the Singular Value Decomposition (SVD) to construct explicit orthogonal semantic and forgery subspaces, avoiding the distortion of well-learned semantic knowledge during fake pattern learning.
@@ 3.2 Effort Approach
where $U_{r}\in\mathbb{R}^{n\times r}$, $\Sigma_{r}\in\mathbb{R}^{r\times r}$, and $V_{r}\in\mathbb{R}^{n\times r}$. We keep $W_{r}$ frozen during training to preserve the knowledge learned from real images.
@@
To encourage the $\Delta W$ to capture both useful and meaningful discrepancy between the real and fake, it’s significant to guarantee that optimizing $\Delta W$ does not change the properties of the overall weight matrix $W$ (i.e., Minimize the impact on the real information of the pre-trained weight as much as possible). Thus, we proposed two constraints to realize this goal:
@@
$\mathcal{L}_{\rm ksv}=\bigg{|}\sum_{i=r+1}^{n}\hat{\sigma}_{i}^{2}-\sum_{i=r+1}^{n}\sigma_{i}^{2}\bigg{|}=\bigg{|}\lVert\hat{W}\rVert_{F}^{2}-\lVert W\rVert_{F}^{2}\bigg{|}$ (6)
@@
where $\lambda_{1},\lambda_{2}$ are hyperparameters that balance the importance of the corresponding regularization term, and $m$ represents the number of pre-trained weight matrices on which our approach is applied.
In practice, we adapt our approach to the linear layers within the self-attention module across all transformer layers of the Vision Foundation Model (VFM) to leverage their rich, well-learned real distributions.

### v2

# Orthogonal Subspace Decomposition for Generalizable AI-Generated Image Detection

@@ Abstract
AI-generated images (AIGIs), such as natural or face images, have become increasingly realistic and indistinguishable, making their detection a critical and pressing challenge. In this paper, we start from a new perspective to excavate the reason behind the failure generalization in AIGI detection, named the asymmetry phenomenon, where a naively trained detector tends to favor overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked, which is proved seriously limiting the expressivity and generalization. One potential remedy is incorporating the pre-trained knowledge within the vision foundation models (higher-ranked) to expand the feature space, alleviating the model’s overfitting to fake. To this end, we employ Singular Value Decomposition (SVD) to decompose the original feature space into two orthogonal subspaces. By freezing the principal components and adapting only the remained components, we preserve the pre-trained knowledge while learning forgery-related patterns. Compared to existing full-parameters and LoRA-based tuning methods, we explicitly ensure orthogonality enabling the higher rank of the whole feature space, effectively minimizing overfitting and enhancing generalization. Extensive experiments with our deep analysis on both deepfake and synthetic image detection benchmarks demonstrate superior generalization performance in detection.
AI-generated Image Detection, Image Decomposition, low-level vision, Deep Learning
Figure 1: Illustration of the asymmetry phenomenon in AI-generated image detection. We show that the baseline detector (i.e., Xception) tends to quickly overfit to the fake patterns in the training set (Rossler et al., 2019), causing the limited generalization when facing previously unseen fakes (Li et al., 2020b).
@@ 1 Introduction
The rapid development of AI generative technologies has significantly lowered the barrier for creating highly realistic fake images. As deep generative models advance and mature (Goodfellow et al., 2020; Ho et al., 2020; Rombach et al., 2022b), the proliferation of AI-generated images (AIGIs [^1] In the context of this research, AIGI primarily refers to deepfakes (face-swapping) and synthetic images (e.g., nature or arts).) has drawn considerable attention, driven by their ability to produce high-quality content with relative ease. However, these advancements also introduce significant risks, if misused for malicious purposes such as deepfakes (mainly including face-swapping (Korshunov & Marcel, 2018) and face-reenactment (Thies et al., 2016)), which may violate personal privacy, spread misinformation, and erode trust in digital media. Consequently, there is an urgent need to develop a reliable and robust framework for detecting AIGIs.
@@
Most existing studies in AIGI detection (Wang et al., 2020b; Rossler et al., 2019) typically approach the real/fake classification problem as a symmetric binary classification task, akin to the “cat versus dog” problem. A standard binary classifier, often based on deep neural networks, is trained to distinguish between real and fake images by predicting the likelihood of a given test image being fake during inference. Although this paradigm yields promising results when the training and testing distributions (in terms of fake generation methods) are similar, its performance tends to degrade significantly when applied to previously unseen fake methods, indicating the generalization issue (Ojha et al., 2023).
To understand the underlying reasons for the failure in generalization, we have conducted extensive preliminary investigations and identified an asymmetry phenomenon in AIGI detection: naively trained detectors tend to take the shortcut and very quickly overfit the limited fake patterns presented in the training set. Visualization in Fig. 1 corroborates this claim. Specifically, the vanilla detector (i.e., Xception (Rossler et al., 2019)) quickly fits the fake patterns at the very early training stage (only a few iterations), resulting in a very low loss of fake, while the real loss is significantly higher than the fake loss ($\sim$ 100$\times$ larger). This is likely because existing AIGI detection datasets (Rossler et al., 2019; Wang et al., 2020b) typically contain limited and homogeneous fake types, while real samples exhibit significantly greater diversity and variance between each other such as different categories and scenarios.
Consequently, the learned feature spaces become fake-dominated and thus highly constrained. As evidenced by the t-SNE visualization in Fig. 2, we see that the whole feature space is indeed dominated by the forgery patterns, where both the Xception (Vanilla CNN) and CLIP (Radford et al., 2021) detector group only the specific fake patterns within the training set into a single cluster, while all other data, including real samples and fake samples from unseen forgeries, are mapped into a separate cluster.
To quantify this, we analyze the effective information [^2] In PCA, it refers to variance captured by principal components. Larger eigenvalues indicate components explaining more variance and contributing more to data representation. contained in the model’s feature space via Principal Component Analysis (PCA). Specifically, we visualize the explained variance ratio of different principal components for the model’s feature space in Fig. 3. The results show that the feature space of the naive detector can be highly constrained and low-ranked, with only two principal components to capture all information, resulting in limited generalization. This aligns with the previous theoretical analysis (Gunasekar et al., 2017) that low-ranked [^3] The “rank” here means the number of significant principal components. feature spaces hinder generalization by memorizing trivial patterns.
Figure 3: Analysis for the effective information contained in the model’s feature space. We apply PCA for dimension reduction and visualize the explained variance ratio of principal components with high contribution. We show that the baseline model trained on the AIGI dataset can be highly constrained and low-ranked.
One potential remedy to address the low-ranked problem is incorporating pre-trained knowledge within vision foundation models (VFMs), which provide higher-ranked feature representations, to expand the low-ranked feature space, thus alleviating the overfitting. However, naively fine-tuning a VFM (even CLIP) risks distorting the original rich representation feature space (Fig. 2), pushing the feature space to become low-ranked again (verified in Fig. 5).
@@
To address this, we design a novel approach called Effort: Efficient orthogonal modeling for generalizable AIGI detection. Specifically, we employ Singular Value Decomposition (SVD) to construct two orthogonal subspaces. By freezing the principal components and adapting the remained components, we preserve the pre-trained knowledge while learning forgery-related patterns.
We have conducted extensive experiments on both deepfake detection and synthetic image detection benchmarks and surprisingly find that our approach achieves significant superiority over other SOTAs with very few training cost. Compared to existing full-parameters and LoRA-based tuning methods, we explicitly ensure orthogonality enabling the higher rank of the whole feature space, effectively minimizing overfitting to fake and enhancing generalization.
Our work makes the following key contributions:
@@
  • Asymmetry phenomenon in AIGI detection: We introduce the concept of asymmetry phenomenon, where a naively trained detector tends to quickly fit the seen fake methods well but, in doing so, it often overfits to specific fake patterns in the training set, limiting its generalization ability to detect unseen fake methods.
  • New perspective to explain the failure reasons behind generalization: We use PCA to quantitatively assess the effective information within the learned feature space, and we find that the overfitting to fake, results in a highly low-ranked and constrained feature space, thus leading to the limited generalization capability.
  • Novel method via orthogonal subspace decomposition: To address the overfitting, we propose a novel approach, Effort, with two careful designs: (1) incorporating the pre-trained knowledge (proving higher-ranked feature space) within the vision foundation models to expand the previous feature space, improving the model’s expressivity and alleviating the overfitting; and (2) employing SVD to explicitly construct two orthogonal subspaces, where the principal one for preserving pre-trained knowledge and the remained one for learning new forgeries, avoiding the distortion of original rich feature space during learning fakes.
@@
  • Stricking generalization performance in detection: Effort achieves remarkably high generalization performance on previously unseen forgeries, which can be applied to detect a broad range of AIGIs, including but not limited to face forgeries (e.g., classical deepfakes) and synthetic natural images.
@@ 2 Related Work
Our work focuses on detecting AI-generated images (AIGIs), especially deepfake images (e.g., face-swapping) and synthetic images (e.g., nature or art), following Yan et al. (2024a). As the majority of recent works specifically focus on dealing with the generalization issue, where the training and testing distribution differ (in terms of fake methods), we will briefly introduce the classical and recent detection methods toward generalization in deepfake and synthetic images, respectively.
@@
The task of deepfake detection grapples profoundly with the issue of generalization. To tackle the generalization issue, one mainstream approach is fake pattern learning. Most existing works are within this line. These methods generally design a “transformation function”, e.g., frequency transformation (Li et al., 2021; Luo et al., 2021; Liu et al., 2021a), blending operations (Li et al., 2020a; Zhao et al., 2021; Shiohara & Yamasaki, 2022; Chen et al., 2022a), reconstruction (Zhu et al., 2021; Cao et al., 2022), content/ID disentanglement (Yan et al., 2023a; Fu et al., 2025; Huang et al., 2023; Dong et al., 2023), to transform the original input $x$ into $x^{\prime}$, where they believe that the more general fake patterns can be captured within the feature space of $x^{\prime}$ compared to $x$. However, given the ever-increasing diversity of forgery methods in the real world, it is unrealistic to elaborate all possible fake patterns and “expect” good generalization on unseen fake methods. Another notable direction is to real distribution learning, with a specific methodology involved: one-class anomaly detection (Khalid & Woo, 2020; Larue et al., 2023). Specifically, Khalid & Woo (2020) introduced a one-class-based anomaly detection, where “abnormal” data is detected by the proposed reconstruction error as the anomaly score. Larue et al. (2023) proposed a similar approach to create pseudo-fake “anomaly” samples by using image-level blending on different facial regions. However, it is challenging to ensure that the detector can learn a robust representation of real images by using the very limited real data in existing deepfake datasets (e.g., the FF++ dataset (Rossler et al., 2019) contains only 1,000 real videos with imbalanced facial attribute distributions (Trinh & Liu, 2021)).
@@
With the rapid advancement of existing AI generative technologies, the scope of forged content has expanded beyond facial forgeries to encompass a wide range of scenes. In this context, similar to the deepfake detection field, most existing works typically focus on fake pattern learning that mines the low-level forgery clues from different aspects. Specifically, several approaches have been proposed to capture low-level artifacts, including RGB data augmentations (Wang et al., 2020b), frequency-based features (Jeong et al., 2022), gradients (Tan et al., 2023), reconstruction artifacts (Wang et al., 2023a; Chen et al., 2024; Luo et al., 2024), and neighboring pixel relationships (Tan et al., 2024c), random-mapping feature (Tan et al., 2024a). To illustrate, BiHPF (Jeong et al., 2022) amplifies artifact magnitudes through the application of dual high-pass filters, while LGrad (Tan et al., 2023) uses gradient information from pre-trained models as artifact representations. NPR (Tan et al., 2024c) introduces a straightforward yet effective artifact representation by rethinking up-sampling operations. In addition to learning from scratch, there are also several research works (Ojha et al., 2023; Wu et al., 2023; Liu et al., 2024) that perform fake pattern learning by leveraging the existing vision foundation models. For instance, UniFD (Ojha et al., 2023) directly freezes the visual encoder of the pre-trained CLIP model and tunes only a linear layer for binary classification, demonstrating effective deepfake detection even with previously unseen sources. LASTED (Wu et al., 2023) proposes designing textual labels to supervise the CLIP vision model through image-text contrastive learning, advancing the field of synthetic image detection. These arts have shown notable improvement in generalization performance when facing previously unseen fake methods.
Figure 4: The proposed approach for AIGI detection. The left branch is the decomposition matrix of the principle components approximation using SVD, while the right residual branch enables the orthogonal learning of real/fake discriminative features.
@@ 3 Methodology
The overall pipeline of the proposed Effort approach is illustrated in Fig. 4, aiming to address the asymmetry phenomena in AIGI detection. Our approach involves the SVD to construct explicit orthogonality for preserving pre-trained knowledge and learning forgery-related patterns, avoiding the distortion of well-learned pre-trained knowledge during learning forgeries.
@@
where $U_{r}\in\mathbb{R}^{n\times r}$, $\Sigma_{r}\in\mathbb{R}^{r\times r}$, and $V_{r}\in\mathbb{R}^{n\times r}$. We keep $W_{r}$ frozen during training to preserve dominant pre-trained knowledge learned from large-scale data.
@@
To encourage the $\Delta W$ to capture both useful and meaningful discrepancy between the real and fake, it’s significant to guarantee that optimizing $\Delta W$ does not change the properties of the overall weight matrix $W$ (i.e., Minimize the impact on the real information of the pre-trained weight as much as possible). Thus, we proposed two constraints to realize this goal, as follows.
@@
$\mathcal{L}_{\rm ksv}=\bigg{|}\sum_{i=r+1}^{n}\hat{\sigma}_{i}^{2}-\sum_{i=r+1}^{n}\sigma_{i}^{2}\bigg{|}=\bigg{|}\lVert\hat{W}\rVert_{F}^{2}-\lVert W\rVert_{F}^{2}\bigg{|},$ (6)
@@
where $\lambda_{1},\lambda_{2}$ are hyperparameters that balance the importance of the corresponding regularization term, and $m$ represents the number of pre-trained weight matrices on which our approach is applied. In practice, we adapt our approach to the linear layers within the self-attention module across all transformer layers of the VFM to leverage their rich, well-learned real distributions.
Theoretical analysis for the asymmetric phenomena in AIGI Detection.
Here, we provide a theoretical explanation for the asymmetric phenomena in AIGI Detection.
Definition 3.1.
(Covariance Asymmetry Measure) Let $\Sigma_{\mathrm{real}}:=\mathbb{E}_{x\sim P_{\mathrm{real}}}[xx^{\top}]$ denote the covariance matrix of real data and $\Sigma_{\mathrm{fake}}:=\mathbb{E}_{z}[\mathbb{E}[x|z]\mathbb{E}[x|z]^{\top}]$ denote the covariance matrix of generator outputs, where $z$ denotes the latent variables. We define the asymmetry metric:
$\gamma:=\frac{\mathrm{rank}(\Sigma_{\mathrm{real}})}{\mathrm{rank}(\Sigma_{\mathrm{fake}})}-1$ (8)
Theorem 3.2.
Under the following assumptions: (i) The real data covariance satisfies $\mathrm{rank}(\Sigma_{\mathrm{real}})\geq r_{\mathrm{real}}$, (ii) Each generator $G_{j}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ satisfies $\mathrm{rank}\left(\mathbb{E}[G_{j}(z)G_{j}(z)^{\top}]\right)\leq r_{\mathrm{fake}}$ for $j=1,\ldots,k$, then with probability exceeding $1-\delta$:
$\gamma\geq\frac{r_{\mathrm{real}}}{kr_{\mathrm{fake}}+C\sqrt{d\log(2k/\delta)}}-1,$ (9)
where $k$ is the number of generators, $d$ is the latent dimension, and $C$ is a universal constant.
This theory establishes the intrinsic asymmetry in AIGI detection through covariance spectral analysis, and uncovers that symmetric classifiers’ inevitable failure as $n/d\rightarrow\infty(\gamma\rightarrow\infty)$, providing mathematical grounding for explaining the observed asymmetric phenomena in AIGI detection. The proof can be seen in the appendix (Sec. A).